OpenStudy (anonymous):

find the product of (x+3) and (x-3)

OpenStudy (anonymous):

use the difference of squares formula a^2-b^2=(a+b)(a-b)

OpenStudy (anonymous):

(x+3)(x-3)=x^2-3^2=.....

OpenStudy (anonymous):

so (x+3)(x-3)=x^2 - 9

OpenStudy (imtiaz7):

we will use formula: different of two square: (a+b)(a-b)=a^2-b^2

OpenStudy (anonymous):

@Laritza does it make sense?

OpenStudy (anonymous):

@Ksanka yes, thank you!